i would like it if you could check my answers on turning point slope form into standard form. the original problem was this: y+5=-5(x-3)i got: 5x+y=10
problem: y-10=-2(x-3) i got: 2x+y=16 problem: y-2=-1/2(x-4) i got: 1/2x+y=4
please check,if they are wrong can you correct. thanks

Answer by lwsshak3(11628)

i would like it if you could check my answers on turning point slope form into standard form. the original problem was this: y+5=-5(x-3)i got: 5x+y=10
problem: y-10=-2(x-3) i got: 2x+y=16 problem: y-2=-1/2(x-4) i got: 1/2x+y=4
**
You got the equations correct but they are not in standard form for straight lines, which is:
y=mx+b, m=slope, b=y-intercept. They should be rewritten in standard form as follows:
..
5x+y=10
y=-5x+10
..
2x+y=16
y=-2x+16
..
1/2x+y=4
y=-x/2+4